# Persuasion in Networks: Can the Sender Do Better than Using Public Signals?

### v2

@@ 4.2 Homophily and type-dependent degree distribution
In this subsection, I study networks that could have homophily and correlation between degrees and types. The conditions for the results in this subsection will be stated in terms of $\{p_{d}^{t}\}_{d}$ and $q_{t}$, i.e. the degree distribution and expected proportion of neighbours of the same type, instead of the distribution over $\lambda$ and $q$. There are two reasons for this. Firstly, from data, it is usually easy to calculate the degree distribution and $q_{t}$, but it is difficult if not impossible to estimate $q$ and the distribution of $\lambda$. Therefore, stating the results in terms of the degree distributions and $q_{t}$ makes it easier to check whether the conditions of the propositions are satisfied. Secondly, all the results would go through when using the more general random networks model from Sadler (2020), which directly imposes arbitrary degree distributions and values of $q_{t}$, so the results are still relevant even we face some degree distribution that could not be generated by the more restrictive model used here. I will also discuss relations between degree distribution, $q_{t}$ and distribution over $\lambda$ and $q$.
I need to introduce some definition to state the results. First, I define forward (degree) distribution. The forward distribution captures the degree distribution of a randomly selected neighbour.
@@
The forward distribution $\{\tilde{p}_{d}^{t}\}_{d}$ for type $t$ is defined as $\tilde{p}_{d}^{t}\equiv\frac{p_{d+1}^{t}(d+1)}{\sum_{d}dp_{d}^{t}}$
Next I define what is meant when I say one type is more connected than the other
@@
I say that type $t$ receivers are weakly more connected than type $t^{\prime}$ if one of the two following conditions hold
@@
Like in the last subsection, we first study the probability that receivers observe viral signals.
@@
Before stating the main proposition, I briefly comment on how the measure $q_{t}$ are related to measures of homophily. Coleman (1958)’s measure of homophily in this model is $\frac{q_{t}-\gamma_{t}}{1-\gamma_{t}}$, and there’s homophily for type $t$ if this measure is positive. Therefore, the network exhibits homophily if $q_{t}>\gamma_{t}$ for both types. Note that if $q_{h}>\gamma_{h}$ and $q_{l}>\gamma_{l}$, then we must have $q_{h}>1-q_{l}$ since $\gamma_{h}$ and $\gamma_{l}$ sum up to 1. Therefore, the condition in the lemma 6 and the proposition below are weaker than requiring the network to exhibit homophily.
Using lemma 6, we can show that conclusions from proposition 1 extend even when the network exhibits homophily, if the believers are weakly more connected than the sceptics.
@@
If the type $h$ receivers are weakly more connected than type $l$ receivers and $q_{h}\geq 1-q_{l}$, conclusions of proposition 1 still hold, i.e.
  1. 1.
    If $v(.)$ is linear or convex, $\limsup_{n\rightarrow\infty}V^{*}_{net}(n)\leq V^{*}_{pub}$.
  2. 2.
    If $\hat{c}>0$, there exists some strictly concave $v(.)$ such that $\liminf_{n\rightarrow\infty}V^{*}_{net}>V^{*}_{pub}$.
Remark 2 also holds so is not stated again for the sake of space.
Homophily is often cited as one of the reasons why advertising and political campaigns on social media is more effective than traditional forms of advertisement, since it influences what information people could observe. Lemma 6 indeed shows that homophily and type-degree correlation introduce asymmetries between the two groups. Roughly speaking, the sender would like to send signals more often to sceptics and keep information away from believers. From Lemma 6, when believers are weakly more connected in the network, the asymmetry introduced is the opposite of what the sender desires.
To discuss the mechanism in more detail, I again impose the simplifying assumption that the sender uses only two signals $s$ and $s^{\prime}$, with $s$ persuading both types and $s^{\prime}$ persuading believers only. From Lemma 6 and focusing on the case where the two types are equally connected, homophily makes believers observe signal $s^{\prime}$ more often than sceptics. This asymmetry means that the imitation argument from last subsection does not directly work. Nonetheless, we could still imitate the probabilities of observing each signals for receivers of a given type and degree $d$. For any degree $d$, there are two cases. If sceptics with degree $d$ choose action 0 upon observing $\emptyset$, then they only choose action 1 when they see $s$. Therefore, if we let them observe signals with the same probabilities as believers with degree $d$, then their probabilities of choosing action 1 will not change. In the second case, suppose sceptics choose action 1 upon observing $\emptyset$, then believers must always choose action 1 when they observe signals with the same probabilities as sceptics since they have higher priors. Therefore, when the sender has convex payoff functions, imitating probabilities for one of the types gives the sender higher payoff than her ’payoff’ from type $d$ receivers (i.e. counterpart here for $V(d,\Pi)$ in last subsection).
@@
Note that this result does not say that homophily is not useful for the sender. Homophily makes believers more clustered together, and signals that only persuade believers can spread more widely. Compared to a network without homophily, sender could send these signals less often while still maintaining the same probability that a believer with degree $d$ observes them. These signals are realised less often now, but are observed by more receivers when realised, which increases risk and potentially makes sender better off.
@@
    If $q=1$, type $t$ is weakly more connected than type $t^{\prime}$ if $f_{t}(\cdot)$ FOSD $f_{t^{\prime}}(\cdot)$ or $f_{t}$ does not depend on $t$.
@@
    Given any $f_{t^{\prime}}(\cdot)$, we can find some $r$ large enough such that type $t$ is more connected than type $t^{\prime}$ if $f_{t}(r\lambda)=f_{t^{\prime}}(\lambda)$ for all $\lambda$.
@@
Finally, I present conditions under which information can be sent to sceptics more often than believers, and therefore network signals are better than public signals even for risk loving sender.
@@
Suppose that there’s no type $l$ receiver with $\lambda$ smaller than some arbitrary $\lambda^{\prime}>0$. We can find some $\overline{x}$ large enough such that if expected degree of type $l$ receivers is larger than $\overline{x}$ and expected degree of type $h$ receivers is smaller than $1/\overline{x}$, then $\liminf_{n\rightarrow\infty}V_{net}^{*}(n)>V_{pub}^{*}$.
@@
The high level intuition underlying this and the last subsection is that, it is difficult to control who sees viral signals using seeding. It is the network structures and how people share signals that determine who observe viral signals. Therefore, whether the sender prefers network signals depends on, roughly speaking, whether the network structure causes believers to receive signals less often than sceptics. From the discussion following proposition 2, it is mainly the probability of observing signals that persuade both types that matters. This is in turn influenced by type-degree correlation. Therefore, whether the sender prefers network signals is mainly determined by whether sceptics are more connected, and homophily cannot reverse the sender’s preference for public signals over network signals when believers are more connected.
@@ 4.3 Voting game
To state the result, I introduce the following definition. Define $d_{vote}$ to be the smallest $d$ such that
$\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}\hat{\zeta}(l,d)}{1-\hat{\zeta}(l,d)}\geq 1$
@@
Roughly speaking, both conditions say that network signals perform better when $\overline{x}-\gamma_{h}$ is small relative to the proportion of type $l$ receivers who could choose action 1 upon observing $\emptyset$. Sender’s payoff function $v$ in voting games is neither concave nor convex, so results in previous sections do not directly apply, but they are still useful to get some intuition for proposition 4. Consider a reduced form problem where sender chooses between always getting moderate levels of support from skeptics, or sometimes getting a lot of support from skeptics and sometimes losing a lot of support from believers. The former option involves less risk. When $\gamma_{h}$ is small and threshold $\overline{x}$ is large, the sender prefers the latter option, since getting moderate support is not enough to reach the threshold. The sender behaves as if she is risk loving, and results in previous subsections suggest that in these cases public signals are better. When $\gamma_{h}$ is close to $\overline{x}$, sender prefers the former option since getting way above the threshold is not valuable. The sender behaves as if she is risk averse, and previous results suggest that network signals can be better.
If skeptics do not choose action 1 upon observing $\emptyset$, they choose action 1 only upon observing signals that persuade both types of receivers. The sender can only get enough support if a large enough fraction of sceptics observe such signals. Therefore, the sender would like to maximise the probability of sending such signals and maximise the proportion of people they reach. This is achieved by using public signals. This explains the importance of having skeptics choosing action 1 upon seeing $\emptyset$ for network signals to be preferred to public signals.
5 Extensions and Discussion
5.1 Alternative sharing rules
@@
As discussed in detail above, the key mechanism behind all the results on public signals being better is that believers observe signals more often than sceptics. These results would still go through with any other rule where (conditional on degree) believers share each signal $s$ more often than sceptics. Such sharing rules ensure that, under conditions stated in the proposition 1 and 2, believers observe each signal more often than sceptics. Note that the sharing rule studied above falls into this more general class, so the proofs are very similar.
@@
I now discuss one sharing rule that does not fall into the class above, but the results in the paper still go through. Suppose that now receivers share a signal $s$ if their action ’agrees’ with content of the signal. Formally, they share the signal if their action is 1 and the signal is more likely in state 1, or if their action is 0 and the signal is more likely in state 0. All receivers will pass on signals that persuade both types and neither types. Only sceptics pass on signals that persuade only the believers, so sceptics can observe these signals more often than believers even if believers are weakly more connected.
@@
However, this cannot make network signals better than public signals if the believers are more connected. To see this, one can go back to the discussion after proposition 2, explaining why public signals are still preferable when network exhibits homophily. Note that the argument there does not use the fact that signals persuading only believers are observed more often by believers. More formally, Lemma 8 in the appendix which presents the formal version of imitation arguments in the main text, does not use that signals persuading only believers are observed more often by believers. Therefore, all results would go through in essentially the same way with this alternative sharing rule.
5.2 Non-vanishing fraction of seeds and optimal network signals
@@
The assumption of seeds being a vanishing fraction of population can be relaxed as long as the seeds only take a fraction of population that is small enough. An upper bound on sender’s payoff can be obtained by assuming that all the seeds and others on the same component with them, excluding those on the giant component, always choose action 1 and remove them from the network. We can then analyze the remaining network and results above will go through. This implies that results would go through as long as the fraction of seeds is small enough even though it is non-vanishing.
@@
For intermediate sizes of seed set, the analysis is more complicated. The main reason is that we now need to consider components other than the giant component. These components are less well studied. Actually, there exist no closed form solutions even for probabilities of being on giant component (i.e. the functions $\zeta$ and $\hat{\zeta}$). Therefore, it is actually difficult to even explicitly determine what receivers do upon observing $\emptyset$ and thus sender’s payoff from a given strategy. This is why, as illustrated above, most results in this paper are obtained from bounding payoffs of network signals from above or below, usually by some imitation argument. Although it would be very interesting to look at sender’s optimal network signal, it is very challenging, since it is not easy to even determine payoff for a given strategy. Introducing intermediate sizes of seed set will make these difficulties more pronounced.
5.3 More than one signal spreading on network
The assumption of having 1 signal spreading on the network is mainly made for tractability. Without this assumption, there may not even exist an equilibrium given some strategies of the sender. To see this, consider an example where there are 2 receivers connected to each other. One of them is a sceptic and the other one is a believer. Suppose that the sender sends a signal that persuades both groups (if it’s the only signal observed) to the sceptic, and a signal that persuades neither to the believer. Also suppose that observing both persuades the believer but not the sceptic. In this case, if sceptic chooses action 1 and shares the signal to the believer, then the believer must also choose action 1. However, this passes on the bad news to the sceptic, which means the sceptic should actually take action 0, and her good news will not be shared in the first place.
One reason this issue arises is that the current model does not explicitly consider time, and assumes that sharing information is simultaneous and instantaneous. This is not an issue with just one signal. With only 1 signal, upon observing some signal, receivers know that they cannot obtain any further information and their posterior will not change. Therefore, we lose no insights by abstracting away from time. However, as illustrated above, we need to introduce time when there are multiple signals. This further introduces the complication of how receivers share information, anticipating that their beliefs may change in the future. The current model is silent on this. This is an interesting direction for future research.

### v3

@@ 4.2 Homophily and type-dependent degree distribution
In this subsection, I study networks that could have homophily and correlation between degrees and types. The conditions for the results in this subsection will be stated in terms of $\{p_{d}^{t}\}_{d}$ and $q_{t}$, i.e. the degree distribution and expected proportion of neighbours of the same type, instead of the distribution over $\lambda$ and $q$. There are two reasons for this. Firstly, from data, it is relatively easy to calculate the degree distribution and $q_{t}$, but it is difficult if not impossible to estimate $q$ and the distribution of $\lambda$. Therefore, stating the results in terms of the degree distributions and $q_{t}$ makes it easier to check whether the conditions of the propositions are satisfied. Secondly, all the results would go through when using the more general random networks model from Sadler (2020), which directly imposes arbitrary degree distributions and values of $q_{t}$, so the results are still relevant even if we face some degree distribution that could not be generated by the more restrictive model used here. I also discuss relations between degree distribution, $q_{t}$ and the distribution over $\lambda$ and $q$ later in this subsection.
I need to introduce some definition to state the results. First, I define the forward (degree) distribution. The forward distribution captures the degree distribution of a randomly selected neighbour.
@@
The forward distribution $\{\tilde{p}_{d}^{t}\}_{d}$ for type $t$ is defined as $\tilde{p}_{d}^{t}\equiv\frac{p_{d+1}^{t}(d+1)}{\sum_{d}dp_{d}^{t}}$ for all $d$.
Next I define what is meant by one type being more connected than the other.
@@
Type $t$ receivers are weakly more connected than type $t^{\prime}$ if one of the two following conditions holds
@@
In words, type $t$ is weakly more connected to type $t^{\prime}$ if they have the same degree distribution, or the degree distribution and forward distribution of type $t$ FOSD those of type $t^{\prime}$ respectively. Like in the last subsection, we first study the probability that receivers observe viral signals.
@@
Before stating the result, I briefly comment on how $q_{t}$ is related to homophily. Coleman (1958)’s measure of homophily in this model is $\frac{q_{t}-\gamma_{t}}{1-\gamma_{t}}$, and there is homophily of type $t$ if this measure is positive. Therefore, the network exhibits homophily if $q_{t}>\gamma_{t}$ for both types. Note that if $q_{h}>\gamma_{h}$ and $q_{l}>\gamma_{l}$, then we must have $q_{h}>1-q_{l}$ since $\gamma_{h}$ and $\gamma_{l}$ sum up to 1. Therefore, the condition in lemma 6 and the proposition below are weaker than requiring the network to exhibit homophily.
Using lemma 6, we can show that conclusions from Proposition 1 extend even when the network exhibits homophily, if believers are weakly more connected than sceptics.
@@
In networks where type $h$ receivers are weakly more connected than type $l$ receivers and $q_{h}\geq 1-q_{l}>0$, a risk-loving or risk neutral sender prefers public signals.
Formally, in such networks, if $v$ is linear or convex, then $\limsup_{n\rightarrow\infty}V^{*}_{net}(n)\leq V^{*}_{pub}$.
It should be emphasized that when setting up the random network model, complete segregation is ruled out by setting $q>0$. This implies that the requirement $1-q_{l}>0$ in the proposition always holds. Note that there can still be very high levels of homophily, i.e. $q_{h}$ and $q_{l}$ can both be arbitrarily close to 1. When there is complete segregation, the sender is obviously in the ideal situation where she could send information only to the sceptics and not to the believers.
Homophily is often cited as one of the reasons why advertising and political campaigns exploiting social networks are more effective than traditional forms of advertisement, because it influences what information people could observe. Lemma 6 indeed shows that homophily and type-degree correlation introduce asymmetries between the two groups. Roughly speaking, the sender would like to send signals more often to sceptics and keep information away from believers. From Lemma 6, when there is homophily and believers are weakly more connected in the network, the asymmetry introduced is the opposite of what the sender desires.
To discuss the mechanism in more detail, I again use the simplified setup for the sketch proof of Proposition 1, where the sender uses only two signals $s$ and $s^{\prime}$, with $s$ persuading both types and $s^{\prime}$ persuading believers only. From Lemma 6 and focussing on the case where the two types are equally connected, homophily makes believers observe signal $s^{\prime}$ more often than sceptics. This asymmetry means that the imitation argument from last subsection does not directly work. Nonetheless, we could still imitate the probabilities of observing each signal for receivers of a given type and degree $d$. For any degree $d$, there are two cases. If sceptics with degree $d$ choose action 0 upon observing $\emptyset$, then they only choose action 1 when they see $s$. Therefore, if we let them observe signals with the same probabilities as believers with degree $d$, then their probabilities of choosing action 1 will not change. In the second case, suppose that sceptics choose action 1 upon observing $\emptyset$, then believers must always choose action 1 when they observe signals with the same probabilities as sceptics since they have higher priors. Therefore, when the sender has weakly convex payoff functions, imitating probabilities for one of the types gives the sender higher payoff than her ’payoff’ from type $d$ receivers (i.e. counterpart here for $V(d,\Pi)$ in last subsection).
@@
This result does not say that homophily is not useful for the sender when she has to use network signals. Homophily makes believers more clustered together, and signals that only persuade believers can spread more widely. Compared to a network without homophily, sender could send these signals less often while still maintaining the same probability that a believer with degree $d$ observes them. These signals are realised less often now, but are observed by more receivers when realised, which increases risk and potentially makes the sender better off.
@@
    If $q=1$, type $t$ is weakly more connected than type $t^{\prime}$ if distribution of $\lambda$ does not depend on type or if $f_{t}$ FOSD $f_{t^{\prime}}$
@@
    Given any $f_{t^{\prime}}$, we can find some $r$ large enough such that type $t$ is more connected than type $t^{\prime}$ if $f_{t}(r\lambda)=f_{t^{\prime}}(\lambda)$ for all $\lambda$.
@@
Next, I present conditions under which information can be sent to sceptics more often than believers, and therefore network signals are better than public signals for the sender.
@@
If expected degree of sceptics is sufficiently large relative to expected degree of believers, the sender prefers network signals.
Formally, suppose that there’s no type $l$ receiver with $\lambda$ smaller than some arbitrary $\lambda^{\prime}>0$. We can find some $\overline{x}$ large enough such that if expected degree of type $l$ receivers is larger than $\overline{x}$ and expected degree of type $h$ receivers is smaller than $1/\overline{x}$, then $\liminf_{n\rightarrow\infty}V_{net}^{*}(n)>V_{pub}^{*}$.
@@
Discussion
One high level intuition underlying this and the last subsection is that, it is difficult to control who sees viral signals using seeding. It is the network structures and how people share signals that determine who observe viral signals. Therefore, whether a risk-loving or risk-neutral sender prefers network signals depends, roughly speaking, on whether the network structure causes believers to receive viral signals less often than sceptics.
As discussed in the literature review, most existing papers on persuasion do not consider endogenous communication between receivers on a network. Communication and the networks impose endogenous restrictions on correlations between what the receivers observe. Lemmas 5 and 6 shed light on these restrictions, and propositions 1 to 3 characterise networks in which public signals can and cannot be outperformed for a risk-loving or risk-neutral sender.
In particular, propositions 1 and 2 show that these restrictions can severely constrain the sender’s payoff using network signals compared to public signals. If the sender can use public signals and knows that network satisfy conditions in these propositions. then more detailed information about the network is not very valuable to the sender. In practice, sending public signals may be more costly than using network signals. However, it is also costly to collect and analyse detailed information about network structure, especially in large and complex networks. Therefore, these results imply that it may be more efficient for the sender to spend resources on sending public signals rather than collecting information about the network.
When sceptics are more connected than believers, Proposition 3 implies that there is a stronger case for understanding and exploiting the network. In this case, the sender has incentive to know how to seed signals in order to make it go viral (i.e. to identify the giant component). Such information allows the sender to get higher payoff than using public signals.
4.3 Risk aversion
Results in the last two subsections are mainly about risk-loving and risk-neutral senders, except for Proposition 3 which puts no condition on the sender’s preference in addition to those in model setup. If the sender is not very risk-averse (i.e. payoff function $v$ is concave but very close to linear), results above still apply. This subsection considers a sufficiently risk-averse sender.
Proposition 4.
If $\hat{c}>0$ (i.e. the giant component in the subnetwork of believers is non-vanishing), there exists a concave payoff function $v$ such that $\liminf_{n\rightarrow\infty}V^{*}_{net}(n)>V^{*}_{pub}$.
In particular [^14] I thank an anonymous referee who suggested this example as a conjecture, suppose that $v(x)=(x^{b}-1)/b$, which has constant relative risk aversion. If $\hat{c}>0$, the sender prefers network signals for sufficiently small $b$.
A necessary and sufficient condition for $\hat{c}>0$ can be found in Bollobás et al. (2007). A sufficient condition is that $\lambda_{min}^{2}\gamma_{h}\sum_{\lambda>0}f_{h}(\lambda)>1$, where $\lambda_{min}$ is the minimum non-zero value $\lambda$ could take. Intuitively, this condition says that believers are sufficiently connected.
With network signals, receivers of the same type can have different observations. When $\hat{c}>0$, by sending signals that persuade only believers to the giant component in the subnetwork of believers, the sender can persuade all believers and at least some sceptics all the time. However, it is in general impossible to persuade all sceptics at the same time using network signals. With public signals, all receivers of the same type must choose the same action, so it is impossible to persuade some sceptics all the time. A sender who is very risk-averse prefers network signals, since it involves less risk.
@@ 4.4 Voting game
To state the result, I introduce the following definition. Define $d_{vote}$ to be
$d_{vote}\equiv\min\{d\in\mathbb{N}_{0}:\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}\hat{\zeta}(l,d)}{1-\hat{\zeta}(l,d)}\geq 1\}$
@@
The following corollary provides conditions on the network that are more intuitive than those in Proposition 5. It highlights the importance of the giant component in the subnetwork of believers. In particular, it provides a condition on the network such that the sender prefers public signals regardless of the threshold $\overline{x}$
Corollary 2.
If $\hat{c}=0$ (i.e. the giant component in the subnetwork of believers is vanishing), then the sender prefers public signals. If $\hat{c}>0$, then the sender prefers network signals if $\overline{x}-\gamma_{h}$ is sufficiently small.
As explained in the last section, whether $\hat{c}$ is positive or not intuitively depends on whether the believers are sufficiently connected in the network.
The intuition and driving forces underlying these results are actually similar to those for Proposition 4 on a risk-averse sender. Both conditions in Proposition 5 require $\overline{x}-\gamma_{h}$ to be small. When this difference is small, the sender is quite risk averse in the sense that her payoff increases sharply by getting a bit more support in addition to that from believers, but does not increase beyond that. The left hand sides for both conditions in Proposition 5 capture the proportion of sceptics that can be made to always choose action 1. Therefore, both conditions in Proposition 5 say that when the sender is very ’risk-averse’, the sender benefits from the ability to always persuade some sceptics using network signals. This is also what drives Proposition 4. Finally, Proposition 4 requires $\hat{c}>0$ for a risk-averse sender to prefer network signals, and the corollary above shows that $\hat{c}=0$ is sufficient for network signals to perform worse in voting games.
When the sender is very risk-averse or when $\overline{x}-\gamma_{h}$ is small, a main factor influencing whether the sender prefers network signals is whether the sender can always persuade all believers and some sceptics. To achieve this, the sender needs to let sceptics observe signals that persuade only the believers. How well the sender can do this in turn depends on whether such signals can get viral among believers (i.e. whether $\hat{c}>0$). This is why $\hat{c}$ is a key measure of the network in this and the last subsection.
Recall that when the sender is risk-loving or risk neutral, what matters is whether sceptics are more connected than believers in the network. We can see that the sender’s preference can have important influence on the aspect of the network that matters for comparison between public signals and network signals.
5 Extensions
Alternative sharing rules:
@@ 5 Extensions
As discussed above in detail, the intuition behind propositions 1 and 2 is that believers observe signals more often than sceptics. These results would still go through with any other rule where (conditional on degree) believers share each signal $s$ more often than sceptics. Such sharing rules ensure that, under the conditions stated in the propositions 1 and 2, believers observe each signal more often than sceptics.
@@
I now discuss one sharing rule that does not fall into this class, but the results still go through. Suppose that now receivers share a signal $s$ if their action agrees with the content of the signal. Formally, they share the signal if their action is 1 and the signal is more likely in state 1, or if their action is 0 and the signal is more likely in state 0. All receivers pass on signals that persuade both types or neither type. Only sceptics pass on signals that persuade only the believers, so sceptics can observe these signals more often than believers even if believers are weakly more connected.
@@
However, this cannot make network signals better than public signals if the believers are more connected. To see this, one can go back to the discussion after proposition 2, explaining why public signals are still preferable when network exhibits homophily. The argument there does not use the fact that signals persuading only believers are observed more often by believers. More formally, Lemma 8 in the appendix, which presents the formal version of imitation arguments in the main text, does not require signals persuading only believers to be observed more often by believers. Therefore, propositions 1 and 2 go through in essentially the same way with this alternative sharing rule.
The driving force for proposition 3 is unchanged by these alternative sharing rules. For the case of a risk-averse sender and voting, the exact network measures that matter can be different from the ones in subsections 4.3 and 4.4. Nonetheless, the high level intuition that the sender cares about always persuading all believers and some sceptics still holds. Therefore, results analogous to those in subsections 4.3 and 4.4 can be obtained.
Public signals do not reach all receivers:
In the analysis above, it is assumed that public signals will reach everyone. In practice, public broadcasts may not reach all the audience.
Suppose that now the realised public signal will be observed by a proportion $y$ of uniformly randomly selected receivers, and the other receivers observe $\emptyset$. Propositions 1 through 3 still hold if $y$ is large relative to size of the giant component. Detailed proofs follow very similar steps, so will be omitted for the sake of space. I illustrate the main idea by discussing when and why the sketch proof of proposition 1 still goes through. Recall that the sketch proof works by using the public signals to imitate probabilities that receivers with degree $d$ observe $s$ and $s^{\prime}$. The public signal should also send $s$ and $s^{\prime}$ with probabilities smaller than those of sending $s$ and $s^{\prime}$ in the network signal structure [^15] This also implies that each of these signals will reach more than $\zeta(d)$ and $\hat{\zeta}(d)$ proportions of receivers with degree $d$. Using network signals, these receivers observe $s$ with probability $\pi(s|\omega)\zeta(d)$ and $s^{\prime}$ with probability $\pi(s^{\prime}|\omega)\hat{\zeta}(d)$. In the current setup, we need the public signal structure to send $s$ and $s^{\prime}$ with probabilities $\pi(s|\omega)\zeta(d)/y$ and $\pi(s^{\prime}|\omega)\hat{\zeta}(d)/y$ respectively. For the imitation argument to work, we need $\zeta(d)/y$ and $\hat{\zeta}(d)/y$ to be less than 1.
Note that $\zeta(d)$ and $\hat{\zeta}(d)$ go to 1 as $d$ goes to infinity, so the imitation argument does not work for high degrees. However, the proportion of nodes with high degrees is small, so their actions do not matter much for the sender’s payoff.
The argument above shows why propositions 1 and 2 still hold for a risk-loving or risk-neutral sender. The intuition for why the sender can prefer network signals when she is risk averse (and $y$ is large enough) or when sceptics are more connected are essentially the same as in the main text.
Non-vanishing fraction of seeds:
@@
The assumption of seeds being a vanishing fraction of population can be relaxed as long as the seeds only take a fraction of population that is small enough. An upper bound on sender’s payoff can be obtained by assuming that all the seeds and others on the same component with them, excluding those on the giant component, always choose action 1 and remove them from the network. We can then analyse the remaining network and the results above will go through. This implies that results would go through as long as the fraction of seeds is small enough even though it is non-vanishing.
@@
For intermediate sizes of seed set, we need to consider components other than the giant component. We need to know the probabilities that a node with type $t$ and degree $d$ is on components of all possible sizes and compositions. As $n$ grows large, the number of probabilities that we need to keep track of for each $t$ and $d$ also goes to infinity. This makes analysis a lot more complicated. It can be shown that, if the sender chooses seeds uniformly randomly, then proposition 1 still goes through. In general, it is difficult to get tractable results.
Sender’s optimal strategy:
Innocenti (2022) characterises optimal public signals with heterogeneous receivers, where the sender’s payoff function is linear. It can be shown that the characterisation still holds for payoff functions used in this paper.
Optimal network signal strategies are very complex mainly for two reasons. First, even though the probabilities that receivers are on giant component (i.e. functions $\zeta$ and $\hat{\zeta}$) can be characterised, there exist no closed-form solutions for them. Secondly, receivers’ optimal actions and thus sender’s payoff are discontinuous at the point where receivers have a posterior of 0.5. Therefore, when receivers’ posterior converge to 0.5 as $n$ gets large, their actions and the sender’s payoff may not converge.
In the online appendix, I show that under certain parameter conditions, the problem of finding the sender’s limiting optimal payoff can be reduced to a problem where the sender is choosing probabilities of two non-empty signals. However, outside these parameters, the sender’s problem can be much more complex. Lemma 2 says that seeds not on the relevant giant component will have little influence. It is also confirmed by the results in section 4 that they do not matter for comparison with public signals. Nevertheless, the online appendix illustrates that these seeds with little influence could be very important for the sender’s optimal payoff using network signals. This is due to the second point in the last paragraph.